# Personalizing story generation by inferring author styles

### v2

@@ 1 Introduction
Figure 1: Our two-stage pipeline for personalized story generation. Stage 1 constructs an Author Writing Sheet with Claim ($C$) and Evidence ($E$) pairs capturing the author’s story-writing characteristics across narrative categories. It is derived from the author’s history of writing prompts ($wp$), author-written stories ($s_{a}$), and LLM-generated Average Stories ($s_{b}$) representing a typical author’s response to the same prompt. Stage 2 uses the Author Writing Sheet to role-play the author, incorporating tailored story rules and a persona description for personalized generation.
Large Language Models (LLMs) are increasingly used in interactive systems to support human users in co-writing tasks (Fang et al., 2024; Yuan et al., 2022; Chung et al., 2022). A key challenge in these settings to generate personalized outputs that align with individual user preferences (Yeh et al., 2024), which improves engagement, agency, and writing quality (Zhao et al., 2025). Personalization is also key in educational applications to deliver effective feedback to students (Anderson et al., 1985), and may support users facing writer’s block (Yuan et al., 2022) or help second language learners in learning a new language (Baffour et al., 2023; Su et al., 2023). Despite these potentials, personalization in story writing has received limited attention.
Story-writing personalization faces several challenges, including a lack of existing resources for evaluating personalization, the inherent subjectivity in user preferences, and methods to evaluate the extent of personalization. To address these gaps, we construct Mythos, a dataset of 3.6k human-written stories ($\approx$1500 tokens each) from 112 authors (16 stories each) across five distinct writing settings: Reddit, AO3, Storium, Narrative Magazine, and New Yorker. We split each author’s stories chronologically into two sets: an earlier Profiling set used to infer their writing characteristics, and a later Generation set used to evaluate how well LLMs can emulate their distinctive writing behavior (Salemi et al., 2024).
@@ 3 Author Writing Sheet
To construct the final Author Writing Sheet $A_{|P|}$, we iteratively merge each intermediate sheet $A_{t}^{\prime}$ with the aggregated sheet $A_{t-1}$, following a moving update approach (Chang et al., 2023). To reduce redundancy and manage output length, we prompt the LLM with a prompt (LLMcombine) to group equivalent Claims, select the best Evidence, and retain ungrouped Claims. We further limit each narrative category to 10 Claim-Evidence pairs prioritizing grouped Claims to fit within the LLM’s context window in subsequent iterations. Each Evidence is tagged with a timestamp indicating its source story.
Similar to Knowledge Tracing in education (Liu et al., 2022; Scarlatos and Lan, 2024), our method processes each story sequentially. Doing so enables efficient updates to the Author Writing Sheet without having to re-process prior stories every time a new story is analyzed (Yeh et al., 2024). Therefore, this method reduces cost and mitigates long-context limitations in LLMs (Li et al., 2024); see Appendix B.1 for a detailed cost analysis.
@@ 6.1 Faithfulness to Writing History
We see that personas have little impact on Sheet, since the overall win-rates for Sheet and Sheet-nP are identical. This result is likely attributed to using the Author Writing Summary, rather than the Sheet, as the reference for faithfulness evaluation. In contrast, for Summ, adding persona descriptions improves win-rates over Summ-nP for amateur sources such as Reddit, AO3, and Storium, but leads to lower performance for expert sources like the N.Mag and N.York. This result is likely due to personas for expert authors failing to capture the nuanced and subtle characteristics of their writing styles, which are reflected in their Author Writing Summary. See Appendix E.1.2 and Tables 11, 12 for examples.
@@ 7.2 Results
Reddit and New Yorker show the highest overall win-rates for the personalized methods (52% and 70%, respectively). Reddit benefits from the broad thematic variety of its crowd-sourced writing prompts (see Figure 3 in the Appendix), which leads to diverse author writing styles and makes authors easier to distinguish from the Average Author (see Table 23). New Yorker, with stories written by experts, benefits from the use of advanced narrative devices such as subtext, which humans can evaluate more reliably than LLMs (Subbiah et al., 2024b) (see Table 24). AO3 and Storium show lower overall win-rates for personalized methods. For AO3, LLMs’ familiarity with common fanfiction tropes [^4] https://archiveofourown.org/admin_posts/25888 strengthens the Average Author baseline (see Table 25). For Storium, the small profile size and weak stylistic differentiation among authors beyond open-ended endings (Xu et al., 2024; Tian et al., 2024) limit the effectiveness of personalization methods (see Table 26). These findings suggest that personalization is most effective when authors explore diverse topics or display distinct traits, such as complex narratives, creativity, or advanced language use. Such traits, reflected in past writings, provide clearer signals for personalizing their style.
@@ 8.1 Personalization and Role-Playing
Recent works have introduced benchmark datasets for personalizing LLM outputs in tasks like email and news writing, focusing on shorter outputs (e.g., 300 tokens for product reviews (Kumar et al., 2024), 850 for news writing (Shashidhar et al., 2024)). These methods infer user traits from history for task-specific personalization (Sun et al., 2024, 2025; Pal et al., 2024; Li et al., 2023; Salemi et al., 2025). In contrast, we address the more subjective task of long-form story writing, with author stories averaging 1500 tokens. Unlike prior role-playing approaches that rely on predefined personas (e.g., Tony Stark, Confucius) (Wang et al., 2024; Sadeq et al., 2024; Tu et al., 2023; Xu et al., 2023), we propose a novel method to infer story-writing personas from an author’s history to guide role-playing.
@@ Appendix A Dataset
In this section, we describe the details of the construction of our dataset which includes five distinct story-writing sources, Reddit [^8] https://www.reddit.com/r/WritingPrompts/, AO3999https://archiveofourown.org/, Storium [^10] https://storium.cs.umass.edu/, Narrative Magazine [^11] https://www.narrativemagazine.com/, and New Yorker [^12] https://www.newyorker.com/.
@@
Dataset Total
Stories
Profile Size Gen Size Total
Authors
Stories/
Author
Token Length
@@
Table 6: Dataset statistics, including the total number of stories, profiling (Profile) and generation (Gen) set distribution, number of authors, average stories per author, and token length in words.
@@ Appendix B Author Writing Sheet
1 Input: profiling set $P=\{(wp_{t},s_{at})\mid t=1,\ldots,|P|\}$,
@@ B.1 Cost Analysis of Author Writing Sheet
In an interactive writing setting, where an author adds a new submission at each time step, the total cost of generating the Author Writing Sheet is 0.06\$ * $|P|$, while the total cost of generating the Author Writing Summary is 0.0015$|P|^{2}$ + 0.019$|P|$. The cost of generating the Author Writing Summary scales quadratically with $|P|$, as the entire author history must be reprocessed at each step, leading to cumulative costs as the profiling set grows. In contrast, the cost of generating the Author Writing Sheet grows linearly, as each new submission incurs a fixed cost of 0.06\$ by running the for loop in Algorithm 1 once.
Figure 12 shows the cost analysis of generating the Author Writing Sheet and Summary using GPT-4o as a function of the profiling set size $|P|$ in an interactive writing setting. Up to $|P|=30$, the cost of generating both remains approximately the same at 1.16\$. However, as $|P|$ increases, the cost of generating the Author Writing Summary grows at a rate of 0.025$|P|$ + 0.30 compared to the Author Writing Sheet which continues to scale linearly. Specifically, for $|P|=100$, the cost of generating the Author Writing Summary is 17\$, compared to 6\$ for the Author Writing Sheet. At $|P|=1000$, the cost of generating the Author Writing Summary escalates to 1520\$, while the cost of generating the Author Writing Sheet remains at 60\$.
@@ F.1 Faithfulness to Author History
Tables 17 and 18 show Faithfulness to Writing History for Llama 3.1 8B and 70B, while Tables 19 and 20 report similarity to the author story. Across both criteria, Oracle achieves higher win-rates with Llama 3.1 70B, indicating that larger models enhance instruction-following and personalization (Chung et al., 2024).

### v3

@@ 1 Introduction
Figure 1: Our two-stage pipeline for personalized story generation. Stage 1 constructs an Author Writing Sheet with Claim ($C$) and Evidence ($E$) pairs capturing the author’s story-writing characteristics across narrative categories. It is derived from the author’s history of writing prompts (wpwp), author-written stories ($s_{a}$), and LLM-generated Average Stories ($s_{b}$) representing a typical author’s response to the same prompt. Stage 2 uses the Author Writing Sheet to role-play the author, incorporating tailored story rules and a persona description for personalized generation.
Large Language Models (LLMs) are increasingly used in interactive systems to support human users in co-writing tasks (Fang et al., 2024; Yuan et al., 2022; Chung et al., 2022). A key challenge in these settings is to generate personalized outputs that align with individual user preferences (Yeh et al., 2024), which improves engagement, agency, and writing quality (Zhao et al., 2025). Personalization is also key in educational applications to deliver effective feedback to students (Anderson et al., 1985; Ashok Kumar and Lan, 2024), and may support users facing writer’s block (Yuan et al., 2022) or help second language learners in learning a new language (Baffour et al., 2023; Su et al., 2023). Despite these potentials, personalization in story writing has received limited attention.
Story-writing personalization faces several challenges, including a lack of existing resources for evaluating personalization and the inherent subjectivity in user preferences. To address these gaps, we construct Mythos, a dataset of 3.6k human-written stories ($\approx$1500 tokens each) from 112 authors (16 stories each [^1] The reported average (16) is the mean of source-level stories-per-author counts (Table 8), while the global average across all authors is 32 because Reddit has more stories per author than New Yorker and Narrative Magazine.) across five distinct writing settings: Reddit, AO3, Storium, Narrative Magazine, and New Yorker. We split each author’s stories chronologically into two sets: an earlier Profiling set used to infer their writing characteristics, and a later Generation set used to evaluate how well LLMs can emulate their distinctive writing behavior (Salemi et al., 2024).
@@ 2 Dataset Collection and Description
Stories per Author Across Sources:
As summarized in Table 8, Reddit authors contribute substantially more stories on average (54 per author), while expert sources such as N.Mag and N.York have only 3–4 stories per author due to limited public availability, reflecting natural differences in data accessibility between amateur and expert sources.
@@ 3 Author Writing Sheet
To construct the final Author Writing Sheet $A_{|P|}$, we iteratively merge each intermediate sheet $A_{t}^{\prime}$ with the aggregated sheet $A_{t-1}$, following a moving update approach (Chang et al., 2024). To reduce redundancy and manage output length, we prompt the LLM with a prompt (LLMcombine) to group equivalent Claims, select the best Evidence, and retain ungrouped Claims. We further limit each narrative category to 10 Claim-Evidence pairs prioritizing grouped Claims to fit within the LLM’s context window in subsequent iterations resulting in a total of 40 Claim-Evidence pairs per author. Each Evidence is tagged with a timestamp indicating its source story.
Similar to Knowledge Tracing in education (Liu et al., 2022; Kumar et al., 2023; Scarlatos et al., 2025), our method processes each story sequentially. Doing so enables efficient updates to the Author Writing Sheet without having to re-process prior stories every time a new story is analyzed (Yeh et al., 2024). Therefore, this method reduces cost and mitigates long-context limitations in LLMs (Li et al., 2024); see Appendix B.1 for a detailed cost analysis.
@@ 6.1 Faithfulness to Writing History
We see that personas have little impact on Sheet, since the overall win-rates for Sheet and Sheet-nP are identical. This result is likely attributed to using the Author Writing Summary, rather than the Sheet, as the reference for faithfulness evaluation. In contrast, for Summ, adding persona descriptions improves win-rates over Summ-nP for amateur sources such as Reddit, AO3, and Storium, but leads to lower performance for expert sources like the N.Mag and N.York. This result is likely due to personas for expert authors failing to capture the nuanced and subtle characteristics of their writing styles, which are reflected in their Author Writing Summary. See Appendix E.1.2 and Tables 13, 14 for examples.
@@ 6.2 Similarity to Author Story
Method k = 1 k = 3 k = 5
Sheet 70 67 66
RAG 38 43 36
Summ 50 56 49
Table 5: Effect of the number of few-shot demonstrations ($k$) for different methods, evaluated on 100 Reddit stories from 20 authors for Similarity to Author Story. The proposed Sheet method outperforms the best of other methods. Best results for each $k$ are bolded.
Sheet performs best across different numbers of few-shot demonstrations:
We analyze the impact of the number of few-shot demonstrations ($k$) on model performance using 100 Reddit stories from 20 authors for different methods. As shown in Table 5, the Sheet method achieves the highest win rate at $k=1$ (70%), with a slight decline as $k$ increases, consistent with observations by (Richardson et al., 2023). In contrast, both RAG and Summ peak at $k=3$ but drop at $k=5$, suggesting sensitivity to prompt length and context dilution. Even at their best settings, these methods lag behind Sheet by 27% and 14%, respectively, demonstrating the superiority of our proposed method.
Method Plot Creat. Dev. Lang. Overall
Sheet (full) 48 74 45 60 70
no Plot 37 65 45 61 49
no Creat. 46 63 51 52 53
no Dev. 37 57 36 46 39
no Lang. 38 61 45 46 44
Table 6: Ablation study on the four narrative categories in the Author Writing Sheet, conducted on 100 stories from 20 Reddit authors for Similarity to Author Story. Each cell reports the win rate (%) of the corresponding method (row) for each narrative category (column). Creat., Dev., and Lang. denote Creativity, Development, and Language Use, respectively. Overall indicates the combined win rate across all categories. Method names such as “No Plot” indicate variants of the Sheet method with the corresponding category excluded. Best numbers for each category are bolded.
All four narrative categories contribute to Sheet’s performance:
To examine the contribution of each narrative category in the Sheet method for similarity to the author story, we conducted an ablation study on a subset of 100 stories from 20 Reddit authors. In this setting, we removed one category at a time from both the story rules and persona descriptions in our Sheet method while keeping the others fixed. As shown in Table 6, we see that, the full model that uses all four categories achieves the highest overall win rate of 70%, outperforming all ablated variants. We also see that, excluding any single category lowers performance on the corresponding narrative dimension, confirming that all four aspects, Plot, Creativity, Development, and Language Use jointly contribute to effective personalization. We further note that removing Development yields the largest drop (31% gap) compared to the full model, indicating that while Development is harder to personalize, it is also essential, as its exclusion results in the lowest overall win rate and highlights its importance to effective personalization.
@@ 7.2 Results
Reddit and New Yorker show the highest overall win-rates for the personalized methods (52% and 70%, respectively). Reddit benefits from the broad thematic variety of its crowd-sourced writing prompts (see Figure 3 in the Appendix), which leads to diverse author writing styles and makes authors easier to distinguish from the Average Author (see Table 25). New Yorker, with stories written by experts, benefits from the use of advanced narrative devices such as subtext, which humans can evaluate more reliably than LLMs (Subbiah et al., 2024b) (see Table 26). AO3 and Storium show lower overall win-rates for personalized methods. For AO3, LLMs’ familiarity with common fanfiction tropes [^5] https://archiveofourown.org/admin_posts/25888 strengthens the Average Author baseline (see Table 27). For Storium, the small profile size and weak stylistic differentiation among authors beyond open-ended endings (Xu et al., 2025; Tian et al., 2024) limit the effectiveness of personalization methods (see Table 28). These findings suggest that personalization is most effective when authors explore diverse topics or display distinct traits, such as complex narratives, creativity, or advanced language use. Such traits, reflected in past writings, provide clearer signals for personalizing their style.
@@ 8.1 Personalization and Role-Playing
Recent works have introduced benchmark datasets for personalizing LLM outputs in tasks like email and news writing, focusing on shorter outputs (e.g., 300 tokens for product reviews (Kumar et al., 2024), 850 for news writing (Shashidhar et al., 2024)). These methods infer user traits from history for task-specific personalization (Sun et al., 2024, 2025; Pal et al., 2025; Li et al., 2023; Salemi et al., 2025). In contrast, we address the more subjective task of long-form story writing, with author stories averaging 1500 tokens. Unlike prior role-playing approaches that rely on predefined personas (e.g., Tony Stark, Confucius) (Wang et al., 2024; Sadeq et al., 2024; Tu et al., 2023; Xu et al., 2023), we propose a novel method to infer story-writing personas from an author’s history to guide role-playing.
@@ 11 Ethical Considerations
Acknowledgments
The authors are partially supported by the NSF under grant IIS-2202506. We thank our Upwork annotators for their dedicated efforts in the annotation process. We are also grateful to Alexander Scarlatos and Jaewook Lee for their valuable discussions and feedback. Finally, we thank members of the UMass ML4Ed and UMass NLP labs for their constructive insights and support on this project.
@@ Appendix A Dataset
In this section, we describe the details of the construction of our dataset which includes five distinct story-writing sources, Reddit [^9] https://www.reddit.com/r/WritingPrompts/, AO3101010https://archiveofourown.org/, Storium [^11] https://storium.cs.umass.edu/, Narrative Magazine [^12] https://www.narrativemagazine.com/, and New Yorker [^13] https://www.newyorker.com/.
@@
Source Total Profile Gen. Authors St./Auth. Tokens
@@
Table 8: Dataset statistics for the five sources in Mythos, reporting total stories (Total), profiling split (Profile), generation split (Gen.), number of authors (Authors), average stories per author (St./Auth.), and average story length in tokens (Tokens).
@@ Appendix B Author Writing Sheet
1Input: profiling set $P=\{(wp_{t},s_{at})\mid t=1,\ldots,|P|\}$,
@@ B.1 Cost Analysis of Author Writing Sheet
[…]
Figure 12 shows the cost analysis of generating the Author Writing Sheet and Summary using GPT-4o as a function of the profiling set size $|P|$ in an interactive writing setting. Up to $|P|=30$, the cost of generating both remains approximately the same at 1.16\$. However, as $|P|$ increases, the cost of generating the Author Writing Summary grows at a rate of 0.025|P||P| + 0.30 compared to the Author Writing Sheet which continues to scale linearly. Specifically, for $|P|=100$, the cost of generating the Author Writing Summary is 17\$, compared to 6\$ for the Author Writing Sheet. At $|P|=1000$, the cost of generating the Author Writing Summary escalates to 1520\$, while the cost of generating the Author Writing Sheet remains at 60\$.
@@ F.1 Faithfulness to Author History
Tables 19 and 20 show Faithfulness to Writing History for Llama 3.1 8B and 70B, while Tables 21 and 22 report similarity to the author story. Across both criteria, Oracle achieves higher win-rates with Llama 3.1 70B, indicating that larger models enhance instruction-following and personalization (Chung et al., 2024).